Operation Summary for March 18, 2011 - March 21, 2011 |
|
The Integrated Luminosity for the period from 3/14/11 to 3/21/11 was 70.75 pb-1 (inverse picobarns). NuMI is off for target replacement. Friday March 18 The day shift began with Operations monitoring the Tevatron (TEV) store 8590 (327.43E30) and the Recycler stash (7.5E10), with the Antiproton Source (Pbar) stacking (49.8E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off for target replacement, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, the Time Line Generator (TLG) stopped responding, which required reverting to an older version of the TLG code to correct the problem. A TEV RF station (TRF3) tripped off. This caused a loss in luminosity of 10E30. Operations established beam to the returning FTBF T-992 experiment located in the Meson area at 2:14 PM. During the evening shift, Controls personnel accessed enclosure B to replace a Meson septa magnet motor. Operations disabled beam to T-992 at 6 PM. Operations terminated store 8590 at 8:10 PM. CDF and D0 each made two-hour accesses to their collision halls. Operators conducted a TEV checkout that lasted through to the end of shift. Saturday March 19 The midnight shift began with Operations monitoring the Recycler stash (378.2E10) and the Pbar stack (7.3E10), and with MiniBooNE taking beam. During the midnight shift, Operations put the TEV into HEP shot setup at 12:25 AM and established store 8593 at 1:27 AM with an average initial luminosity (AIL) of 347.12E30. Operations reported beam position problems in the Pbar AP1 line. Operators as well as personnel from Pbar, Main Injector, and Booster looked at the beam positions throughout the shift. Operations reported that the Booster chopper started sparking and caused two hours of downtime. Operators called in Booster personnel. Operations established beam to T-992 at 6:22 AM. During the day shift, TRF1 tripped off and caused a drop in luminosity of 17E30. Operators and other personnel continued to tune on the beam positions for the Pbar AP1 line and had them corrected by noon. During the evening shift, Operations disabled beam to T-992 at 6 PM. Operations reported that Booster and Linac RF systems had caused minor downtime. Operations terminated store 8593 at 8:53 PM, put the TEV into shot setup, and established store 8594 at 10:06 PM with an AIL of 391.6E30. Sunday March 20 The midnight shift began with Operations monitoring store 8594 and the stash, and with Pbar stacking, and with MiniBooNE taking beam. During the midnight shift, Operations established beam to T-992 at 6:56 AM. Operations reported minor downtimes caused by Linac RF systems. During the day shift, at 10:48 AM, several lightning strikes caused several problems: store 8594 aborted, the stochastic cooling for Recycler was lost as well as part of the stash, part of the Pbar Accumulator stochastic cooling system was lost, and an interlocked beam detector at MI-12 failed. Operators recovered stacking and reestablished beam to T-992 at 10:55AM. Interlock technicians repaired the MI-12 beam detector hardware. Operations restored beam to MiniBooNE at 12:22 PM. During the evening shift, Operations disabled beam to T-992 at 6 PM. Operators found and replaced three bad attenuator modules at 7 PM that were causing the Recycler cooling problems. Operations transferred antiprotons from the Accumulator to Recycler in preparation for a Pbar rings access. Operations reported multiple issues turning the Tevatron back on. The TEV was quenching repeatedly at sector D17 at an energy of 90 GeV. Operators checked the cryo system and changed a VFC card, but to no avail. Operators accessed the TEV to verify the status of the turbo vacuum pumps. Operators prepared the Pbar rings for access at 8:26 PM. Pbar personnel replaced a laser receiver for the 4-8 momentum system along with a 12-volt power supply in the upstairs PLC rack. Pbar resumed stacking at 10:53 PM. Operations attempted to ramp the TEV to full field; unfortunately, it quenched at D17 once more. TEV personnel reported that they were able to turn on to a 90 GeV null ramp and then transition to a 200 GeV ramp. At 11:55 PM, operators prepared the TEV for access to install a portable quench detection (PDQ) monitor at D17. Monday March 21 The midnight shift began with the TEV in access, with Operations monitoring the stash, and with Pbar stacking, and with MiniBooNE taking beam. During the midnight shift, TEV personnel reported at 5:09 AM that they had completed installing the PDQ monitor. Operations began the TEV turn on and TEV personnel began their quench investigation with the PDQ. Operations established beam to T-992 at 4:26 AM. The Plans for Monday The plans for today are to prepare for a sector D TEV magnet (D16-2) replacement. This will require warming up sector D. The magnet replacement and cryo system recovery will take 10-11 days. |
|
More Information Comments and Suggestions |
| last modified 03/21/2011 email Fermilab |
| Security, Privacy, Legal | |